Circular No. 6621

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE The Supernova Cosmology Project (cf. IAUC 6540, plus A. Fisher)

report nine supernovae found with the Cerro Tololo (CTIO) 4-m telescope in their Spring 1996 High Redshift Supernova Search:

SN  1996 UT  α_{2000}  δ_{2000}  R  z  type T
1996cf Mar. 17 10^{h}48^{m}50.^{s}96 + 0^{o}03′30′′.4 22.7 0.57* Ia  M
1996cg Mar. 17 8 24 13.33 + 3 24 58.5 22.5 0.47* Ia K,M
1996ch Mar. 17 11 59 37.73 − 9 54 34.2 24.1 0.57 Ia  K
1996ci Mar. 17 13 45 56.16 + 2 26 53.7 22.3 0.50 Ia K,M
1996cj Mar. 17 13 47 37.81 + 2 26 14.9 23.5 0.19 II  K,E
1996ck Mar. 17 12 48 35.19 − 0 46 03.0 23.0 0.66 Ia? K
1996cl Mar. 18 10 56 59.13 − 3 37 36.4 24.0 0.83 Ia? K,M
1996cm Mar. 18 15 30 11.25 + 5 55 27.0 22.7 0.45* Ia E,K
1996cn Mar. 18 13 48 27.22 + 2 27 46.6 22.6 0.44* Ia  E

The supernovae were all discovered before or at maximum light; they were reobserved during March–May at CTIO, WIYN, and the Isaac Newton Telescope, and the spectra were obtained at the Keck (K) telescope on Apr. 25 and May 12, at the Multiple Mirror Telescope (M) on Mar. 19–20 by S. Pascarelle and R. Windhorst, and at the ESO 3.6-m (E) telescope on Mar. 23 and 24 by N. Caon and S. Benetti and on Mar. 27 by D. Koester. The last column in the table indicates the telescope for the spectroscopic observation. The redshifts were determined from host-galaxy emission and absorption lines, except for those marked with an asterisk, which were determined solely from the spectra of the supernova. The supernovae are all within 3′′.0 of the host-galaxy’s center.

Corrigendum. On IAUC 6596, line 2, for (INT)] read (INT), and S. Fabbro at Berkeley Laboratories]”

SUPERNOVA 1997bq IN NGC 3147 S. Nakano, Sumoto, Japan, forwards the following precise position of

SN1997bq (cf. IAUC 6616), obtained by M. Aoki, Toyama, on Apr. 9.628 UT with a 0.43-m f/6 reflector (+ unfiltered CCD): α = 10^{h}17^{m}05^{s}.33, δ = +73^{o}23′02′′.1 (equinox 2000.0), which is 50′′ east and 60′′ south of the center of NGC 3147. The supernova’s magnitude was 15.4.

1997 April 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green